the sum of claire's and hezron's ages are 50 years. in 2 1/2 years claire will be 5/6 of hezron's age. find there present ages
Answer by sachi(548)

let claire's age is x
and hezron's age is 50-x
in 2 1/2 yrs
claire's age will be x+2.5
and hezron's age will be 50-x+2.5=52.5-x
as per question
x+2.5=5/6(52.5-x)
or 6x+15=262.5-5x
or 11x=262.5-15=247.5
or x= 22.5 yrs
so claire's age is 22.5 yrs
and hezron's age is 50-22.5=27.5 yrs ans